OpenStudy (anonymous):

How do you solve 63 ----- 8 with a radical sign over the whole thing?

OpenStudy (anonymous):

\[\sqrt{63/8}\]

OpenStudy (anonymous):

are you trying to rationalize the denominator?

OpenStudy (anonymous):

simplifying the whole thing.

OpenStudy (anonymous):

\[63=9\times 7\] so \[\sqrt{36}=\sqrt{9}\times \sqrt{7}=3\sqrt{7}\] and similarly \[\sqrt{8}=2\sqrt{2}\]so start with \[\frac{3\sqrt{7}}{2\sqrt{2}}\]

OpenStudy (anonymous):

then you can multiply top and bottom by \[\sqrt{2}\] if you like to get \[\frac{3\sqrt{7}}{2\sqrt{2}}\times \frac{\sqrt{2}}{\sqrt{2}}=\frac{3\sqrt{14}}{4}\]\]